OpenStudy (anonymous):

YES OR NO? The lengths of the sides of a triangle are 6, 8, 10. Can the triangle be a right triangle?

myininaya (myininaya):

\[6^2+8^2=? 10^2\]

OpenStudy (hoblos):

6^2 + 8^2 = 36+64 = 100 = 10^2 ==> yes